# Compact 3D Scene Representation via Self-Organizing Gaussian Grids

### v1

@@ Abstract
3D Gaussian Splatting has recently emerged as a highly promising technique for modeling of static 3D scenes. In contrast to Neural Radiance Fields, it utilizes efficient rasterization allowing for very fast rendering at high-quality. However, the storage size is significantly higher, which hinders practical deployment, e.g. on resource constrained devices. In this paper, we introduce a compact scene representation organizing the parameters of 3D Gaussian Splatting (3DGS) into a 2D grid with local homogeneity, ensuring a drastic reduction in storage requirements without compromising visual quality during rendering. Central to our idea is the explicit exploitation of perceptual redundancies present in natural scenes. In essence, the inherent nature of a scene allows for numerous permutations of Gaussian parameters to equivalently represent it. To this end, we propose a novel highly parallel algorithm that regularly arranges the high-dimensional Gaussian parameters into a 2D grid while preserving their neighborhood structure. During training, we further enforce local smoothness between the sorted parameters in the grid. The uncompressed Gaussians use the same structure as 3DGS, ensuring a seamless integration with established renderers. Our method achieves a reduction factor of 8x to 26x in size for complex scenes with no increase in training time, marking a substantial leap forward in the domain of 3D scene distribution and consumption. Additional information can be found on our project page:
fraunhoferhhi.github.io/Self-Organizing-Gaussians/
Figure 1: Our 3DGS training method allows for high compression of 3D Gaussian Splatting attributes while maintaining high rendering quality. By sorting the Gaussian features periodically into a 2D grid and applying a smoothness loss during scene generation, we are able to significantly reduce the memory footprint using state-of-the-art image compression methods.
@@ 1 Introduction
To address this limitation, we propose a new compact representation based on 3DGS, allowing for efficient encoding and storage. We exploit the fact that numerous permutations of Gaussian parameters can equivalently represent it with the same visual quality by arranging the millions of multidimensional list entries of 3DGS into a into a structured 2D grid with increased smoothness that can efficiently be encoded. In essence, we integrate a smoothness loss into the 3DGS scene optimization process to guide the ambiguous arrangements of the splats towards a visually equivalent but better compressible parameter set. We finally quantize the parameters and encode them by efficient 2D image compression methods, reducing the overall memory footprint by a factor of 8x to 26x, without sacrificing rendering quality.
With our method, we enable novel-view synthesis at high quality, high rendering speed, and with a small memory footprint, bringing high quality 3D scene reconstruction and rendering one step closer towards applications on small devices with limited memory or fast web applications.
@@
    We introduce an efficient 2D sorting algorithm that sorts millions of 3DGS parameters in parallel on the GPU.
@@
    We efficiently reduce the storage size by a factor of 8x to 26x while maintaining high visual quality
@@ 2 State of the Art
Early novel-view synthesis approaches were based on pure image interpolation as for the lightfields [10, 19], or reconstructed 3D geometry by means of regular voxel grids [29], point clouds, or meshes. Structure-from-Motion (SfM) and Multi-View-Stereo (MVS) were proposed to estimate such representations from a set of input images [5, 31, 9]. These methods were the basis for view-synthesis approaches guided by explicit representations of the geometry [3, 17, 28]. Also compression of the large 3D point cloud datasets has been addressed [11, 26], but is typically restricted to simple colored points without additional attributes.
@@
The success of NeRF has resulted in numerous follow-up methods improving rendering quality, e.g. for large scenes [23], or speed [1, 25, 7, 35]. While Mip-NeRF360 [1] as one of the state-of-the-art approaches achieves an outstanding quality, it is still very costly to train and render. Methods focusing on faster training and/or rendering exploit spatial data structures and encodings in order to accelerate computation, representing the scene as a grid [7], tensors [4], latent features [21, 13, 34], hash grids [25] or using vector quantizations [35]. Among these, InstantNGP is a widely recognized example, using a smaller MLP to represent density and appearance though the use of a hash table and an occupancy grid [25]. Plenoxels estimate explicit spherical harmonics in addition to color and density to represent view dependent appearance effects [7]. In order to improve the rendering, the values are interpolated from a sparse 3D grid of features using trilinear interpolation, resulting in considerably higher rendering speed at the cost of lower quality. In addition, the memory size of Plenoxels is high, given that the number of elements in the 3D grid grows by $\mathcal{O}(n^{3})$.
@@
Our approach is inspired by a recently proposed approach for 3D scene representation by Kerbl et al. [14] combining the advantages of explicit representations (fast training and rendering) with those of implicit representation (stochastic sampling) by representing the scene by millions of 3D Gaussians, called 3D Gaussian Splatting (3DGS). The parameters (3D coordinates, color, size, density, orientation, and spherical harmonics) of these Gaussians are optimized using an efficient differentiable rasterization algorithm starting from a set of Structure-from-Motion points to best represent a given 3D scene. The main contribution of this approach is a representation that is continuous and differentiable while also allowing fast rendering: 3DGS is currently the state-of-the-art approach in terms of quality, rendering time, and performance. However, this comes at the cost of significantly higher memory requirements, as for each scene, millions of high dimensional Gaussians are stored in an unstructured list.
@@ 2.2 Mapping Higher Dimensional Data to 2D Grids
A self-organizing map (SOM) [15, 16] is a neural network trained using unsupervised learning to produce a low-dimensional (typically two-dimensional), discretized representation of the input space of the training samples. Initialized with random vectors, the map is refined through iterative adjustments to represent the input data in a topologically organized manner until convergence. Due to the sequential nature of the SOM, the last input vectors can only be assigned to the few remaining unassigned map positions, resulting in isolated and poorly positioned vectors within the map. Self-Sorting Maps [33, 32] avoid this problem by swapping assigned positions of four input vectors at a time. Due to the factorial number of permutations, adding more candidates would be computationally too complex. Linear Assignment Sorting (LAS) [2] outperforms other sorting methods in terms of speed and sorting quality for non-parallel computations, by swapping all (or several (Fast LAS)) vectors simultaneously using a continuously filtered map.
@@
3 Method
Our goal is a new representation and training scheme for 3D Gaussian Splatting with drastically reduced storage requirements while maintaining visual quality during rendering. The idea of our method is to map originally unstructured 3D Gaussians to a 2D grid in such a way that local spatial relationships are preserved. By mapping all Gaussian parameters (position, color, etc.) to the same 2D position, we effectively create multiple data layers with the same layout. In this 2-dimensional layout, neighboring Gaussians will have similar attribute values, facilitating compression and enabling the storage of the same scene using reduced memory.
Figure 2 illustrates an outline of our method. Our training introduces two new components to the 3DGS framework: a highly-parallel 2D sorting step (Section 3.1), and an enhanced scene optimization with a local neighborhood loss on the 2D grid (Section 3.2). The sorting step initializes the optimization process by explicitly organizing the values, establishing an initial smooth local neighborhood without requiring a global optimization. Once sorted, our enhanced training incorporates a neighborhood loss, further enforcing local smoothness on the 2D grid, while the differentiable renderer of the 3D Gaussian splatting optimizes the Gaussians attributes to accurately represent the scene. Similar to the original 3DGS approach, Gaussians are pruned from the scene in regular intervals, and new Gaussians are added in under-reconstructed parts of the scene. After each densification step, we re-sort all Gaussians into the 2D grid. This step re-establishes local neighborhoods after pruning and densification may have disrupted them.
To store the Gaussian attributes, we leverage off-the-shelf compression methods, which can be applied effectively to the well-organized data. Inspired by [1], we map the position of the Gaussians to a logarithmic scale. Thus, Gaussians located in the center of the scene have more precise coordinates.
@@ 3.1 Sorting High Dimensional Gaussians into 2D Grids
Traditional grid-sorting algorithms are suitable for thousands of items, yet fall short for our application, where even small scenes contain millions of Gaussians across multiple dimensions. To address this, we combine and enhance concepts from existing sorting methods discussed in Section 2.2. Our aim is to handle the increased data volume and complexity efficiently, while ensuring that regular sorting of Gaussian data during training does not negatively affect training time.
Our strategy involves an approximation approach to the highly complex assignment problem, inspired by the principles of the Fast Linear Assignment Sorting (FLAS) algorithm [2]. We aim to find an optimal balance between computational efficiency and maintaining the accuracy of the spatial relationships among the Gaussian attributes. The parallel assignment process is visualized with random RGB colors in Figure 3: We initialize our grids by mapping the Gaussians to random positions, even if they are already sorted, to avoid getting stuck in local minima. Subsequently, we perform a low-pass filtering operation on the grid to construct an idealized target grid, and re-assign all elements to their best-matching positions regarding the smooth target. This process is repeated while decreasing the filter size, gradually sorting the grid.
@@
Organizing Gaussians into a 2D Grid: The presented algorithm can be implemented highly parallel on the GPU, and will organize millions of Gaussians in a few seconds, using a few thousand iterations of the blockwise linear assignment. In our application, we normalize all attributes of the Gaussians (coordinates, scale, etc.) independently to the range $[0,1]$ and then scale them with individual weights before calculating the distance of the possible re-assignments. We use the activated values for all attributes (e.g. exponentially activated scale), as they would be seen by the renderer. At the start of the training, we sort once, and continue sorting with each densification step. As the number of Gaussians grows over time, so will the the size of the grid. We find the largest square grid size that can be completely filled with Gaussians for simplicity, and prune the Gaussians with the lowest opacity values that will not fit.
@@ 3.2 Smoothness Regularization
For the regularization, we first create a blurred and thus smoother version of the 2D grid by applying a 2D Gaussian filter on all attribute channels. Deviations of the current Gaussian parameters at an original 2D grid position from those of the smoothed grid form the additional smoothness loss for training. For comparison, we select a Huber loss, as it is less sensitive to outliers than MSE. This has shown to improve the compression rate. Formally, our smoothness regularization term can be expressed as follows:
$R_{s}(I)=\begin{cases}\frac{1}{2}(I-I_{b})^{2},&\text{for }|I-I_{b}|\leq 1\\ |I-I_{b}|-\frac{1}{2},&\text{otherwise.}\end{cases}$ (1)
Here, $I$ and $I_{b}$ denote the original and blurred 2D attribute grid, respectively. With this formulation, the smoothness regularization penalizes all pixels on the 2D grid that have very different values compared to their local neighborhood, thus creating smoother 2D features. We add the loss of the smoothness regularization to the loss of the 3D Gaussian Splatting algorithm using a weight $\lambda$. As we propagate the gradient of the smoothness regularization through the Gaussian renderer during training, we specifically enforce the 3D Gaussian Splatting optimization process to prefer Gaussians that improve the quality of the rendering while also considering the respective local smoothness. This differentiates our method from compression methods that focus on reducing existing data in a pure post-processing step. Instead, we manipulate the data during creation to achieve high compression and high visual quality afterwards.
@@
Table 1: A comparison of our method to the default 3DGS with and without spherical harmonics (SH) and prior NeRF-based renderer. Compared to 3DGS without spherical harmonics, a size reduction between 8.8x and 26.5x is reached, depending on the dataset. All sizes are in MB. Results with a $\dagger$ are directly copied from [14, 20]. Results for SSIM are included in the supplementary material.
4 Evaluation
4.1 Implementation Details
Our experiments are based off the official 3DGS implementation on GitHub [^1] https://github.com/graphdeco-inria/gaussian-splatting. The parallel sorting is implemented in PyTorch and also uses the CUDA backend. For our comparisons to prior methods, we select three real-world 3D scene reconstruction datasets: Mip-NeRF360, Tanks&Temples and Deep Blending and a synthetic dataset: Synthetic-NeRF.
During optimization, we have found two sets of parameters that showed promising results in terms of rendering quality and size. In the following we will refer to those configurations as Q and S. Q denotes the configuration that mainly focuses on rendering quality, while S rather focuses on size. A more detailed description of both parameter configurations is given in the supplementary material of this paper. For the compression of our 2D attribute grids, we use a combination of EXR and JPEG-XL. A more comprehensive study on different compression methods in combination with different quantization levels is also given in the supplementary material. For all our experiments, we disable the spherical harmonics from 3DGS.
4.2 Results
@@
Quantitative results: In table 1 we compare our method to the default 3D Gaussian Splatting algorithm and prior NeRF-based 3D reconstruction methods, reporting the standard PSNR, L-PIPS, and SSIM metrics. We demonstrate that we reduce the average memory size by a factor of 8x to 26x compared to 3DGS [14] without spherical harmonics, without sacrificing visual quality. The highest reduction is observed on the Deep Blending dataset. For some datasets, compared to Mip-NeRF360 [1] and VQ-TensoRF [20], we achieve a slightly higher PSNR and slightly lower L-PIPS, but at the same time allow for real time rendering and fast training, given the efficient rasterization algorithm of 3DGS [14]. Additionally, while Mip-NeRF360 trains one scene for several hours, our method only uses 10 to 30 minutes. Most notably, we achieve the same training time as 3DGS, even though our method brings some computational overhead compared to the default 3DGS, arising from the periodic sorting of Gaussians and the necessity to blur grids for computing the neighbor loss. This is due to our parameter configuration enforcing the optimization process to create a considerably smaller number of Gaussians. For example, for the Truck dataset, our method creates 1.55M Gaussians, whereas the default 3DGS creates 2.58M.
For the same reason of the reduced number of Gaussians, our scenes render much faster in the 3DGS viewer than the original models. The vanilla Truck with 2.58M Gaussians renders at 385 fps with the default settings on the same GPU, while our scene with 1.55M Gaussians renders at 515 fps with nearly the same visual quality.
@@
Qualitative results: The results of the metrics suggest that the rendering quality of our method is very similar to the vanilla 3DGS. This is also demonstrated in figure 5, where we directly compare the renderings for the truck scene side by side. Here, the training with activated spherical harmonics shows the best rendering results, especially when inspecting transparent materials such as the windshield of the truck. Besides that, all other renderings that do not use spherical harmonics show very similar quality. This underlines that our method is able to maintain a high rendering quality, while only using a fraction of the memory.
Figure 5: A qualitative comparison on the truck dataset between our method and 3DGS with and without spherical harmonics. Ours-Q denotes the configuration that focuses on rendering quality, while Ours-S focuses on disk size.
@@
Figure 6 displays examples of the sorted 2D grids during the training process. Specifically, we show the 2D RGB color grids for the counter, the flowers and the kitchen dataset. We observe that the colors of the scenes organize themselves into clusters of similar colors. The grids, however, do not appear to be more organized with more iterations. This has two reasons. Firstly, the number of Gaussians grows substantially during the training, posing an increasingly more difficult sorting task. And secondly, as our method has to sort all attributes at once using the same permutation, isolated features, such as the color can not be organized perfectly.
@@
Figure 6: Resulting 2D attribute grids for the colors for the respective scene during training. Until iteration 15000, the number of Gaussians grow from about 100k-200k up to 1.5-2 million.

### v2

@@ Abstract
3D Gaussian Splatting has recently emerged as a highly promising technique for modeling of static 3D scenes. In contrast to Neural Radiance Fields, it utilizes efficient rasterization allowing for very fast rendering at high-quality. However, the storage size is significantly higher, which hinders practical deployment, e.g. on resource constrained devices. In this paper, we introduce a compact scene representation organizing the parameters of 3D Gaussian Splatting (3DGS) into a 2D grid with local homogeneity, ensuring a drastic reduction in storage requirements without compromising visual quality during rendering. Central to our idea is the explicit exploitation of perceptual redundancies present in natural scenes. In essence, the inherent nature of a scene allows for numerous permutations of Gaussian parameters to equivalently represent it. To this end, we propose a novel highly parallel algorithm that regularly arranges the high-dimensional Gaussian parameters into a 2D grid while preserving their neighborhood structure. During training, we further enforce local smoothness between the sorted parameters in the grid. The uncompressed Gaussians use the same structure as 3DGS, ensuring a seamless integration with established renderers. Our method achieves a reduction factor of 17x to 42x in size for complex scenes with no increase in training time, marking a substantial leap forward in the domain of 3D scene distribution and consumption. Additional information can be found on our project page: fraunhoferhhi.github.io/Self-Organizing-Gaussians/
Figure 1: Our 3DGS training method allows for high compression of 3DGS attributes while maintaining high rendering quality. By sorting the Gaussian features periodically into a 2D grid and applying a smoothness loss during scene generation, we significantly reduce the storage footprint using state-of-the-art image compression methods.
@@ 1 Introduction
To address this limitation, we propose a new compact representation based on 3DGS, allowing for efficient encoding and storage. We exploit the fact that numerous permutations of Gaussian parameters can equivalently represent a scene with the same visual quality and arrange the millions of multidimensional list entries of 3DGS into a structured 2D grid with increased smoothness for efficientencoding.
In addition, we integrate a smoothness loss into the 3DGS scene optimization process to guide the ambiguous arrangements of the splats towards a visually equivalent but better compressible parameter set. We finally quantize the parameters and encode them by efficient 2D image compression methods, reducing the overall file size by a factor of 17x to 42x, without sacrificing rendering quality.
With our method, we enable novel-view synthesis at high quality, high rendering speed, and with a small storage footprint, bringing high quality 3D scene reconstruction and rendering one step closer towards applications on small devices with limited storage capacity or fast web applications.
@@
    We introduce an efficient 2D sorting algorithm called Parallel Linear Assignment Sorting (PLAS) that sorts millions of 3DGS parameters on the GPU in seconds.
@@
    We efficiently reduce the storage size by a factor of 17x to 42x while maintaining high visual quality
@@ 2 State of the Art
Early novel-view synthesis approaches were based on pure image interpolation as for the lightfields [11, 20], or reconstructed 3D geometry by means of regular voxel grids [32], point clouds, or meshes. Structure-from-Motion (SfM) and Multi-View-Stereo (MVS) were proposed to estimate such representations from a set of input images [5, 34, 9]. These methods were the basis for view-synthesis approaches guided by explicit representations of the geometry [3, 18, 31]. Also, compression of the large 3D point cloud datasets has been addressed [12, 29], but is typically restricted to simple colored points without additional attributes.
@@
The success of NeRF has resulted in numerous follow-up methods improving rendering quality, e.g. for large scenes [25], or speed [1, 27, 7, 38]. While Mip-NeRF360 [1] as one of the state-of-the-art approaches achieves an outstanding quality, it is still very costly to train and render. Methods focusing on faster training and/or rendering exploit spatial data structures and encodings in order to accelerate computation, representing the scene as a grid [7], tensors [4], latent features [23, 14, 37], hash grids [27] or using vector quantizations [38]. Among these, InstantNGP is a widely recognized example, using a smaller MLP to represent density and appearance though the use of a hash table and an occupancy grid [27]. Plenoxels estimate explicit spherical harmonics in addition to color and density to represent view dependent appearance effects [7]. In order to improve the rendering, the values are interpolated from a sparse 3D grid of features using trilinear interpolation, resulting in considerably higher rendering speed at the cost of lower quality. In addition, the storage size of Plenoxels is high, given that the number of elements in the 3D grid grows by $\mathcal{O}(n^{3})$.
@@
Our approach is inspired by a recently proposed approach for 3D scene representation by Kerbl et al. [15] combining the advantages of explicit representations (fast training and rendering) with those of implicit representation (stochastic sampling) by representing the scene by millions of 3D Gaussians, called 3D Gaussian Splatting (3DGS). The parameters (3D coordinates, color, size, density, orientation, and spherical harmonics) of these Gaussians are optimized using an efficient differentiable rasterization algorithm starting from a set of Structure-from-Motion points to best represent a given 3D scene. The main contribution of this approach is a representation that is continuous and differentiable while also allowing fast rendering: 3DGS is currently the state-of-the-art approach in terms of quality, rendering time, and performance. However, this comes at the cost of significantly higher storage requirements, as for each scene, millions of high dimensional Gaussians are stored in an unstructured list on disk.
@@ 2.2 Mapping Higher Dimensional Data to 2D Grids
A self-organizing map (SOM) [16, 17] is a neural network trained using unsupervised learning to produce a low-dimensional (typically two-dimensional), discretized representation of the input space of the training samples. Due to the sequential nature of the SOM training algorithm, the last input vectors can only be assigned to the few remaining unassigned map positions, resulting in isolated and poorly positioned vectors within the map. Self-Sorting Maps [36, 35] avoid this problem by swapping assigned positions of four input vectors at a time. Due to the factorial number of permutations, adding more candidates would be computationally too complex. Linear Assignment Sorting (LAS) [2] outperforms other sorting methods in terms of speed and sorting quality for non-parallel computations, by swapping all (or several (Fast LAS)) vectors simultaneously using a continuously filtered map.
@@
Our goal is a new representation and training scheme for 3D Gaussian Splatting with drastically reduced storage requirements while maintaining visual quality during rendering. The idea of our method is to map originally unstructured 3D Gaussians to a 2D grid in such a way that local spatial relationships are preserved. By mapping all Gaussian parameters (position, color, etc.) to the same 2D position, we effectively create multiple data layers with the same layout. In this 2-dimensional layout, neighboring Gaussians will have similar attribute values, facilitating compression and enabling the storage of the same scene with much less disk space.
3 Method
Figure 2 illustrates an outline of our method. Our training introduces two new components to the 3DGS framework: a highly-parallel 2D sorting step (Section 3.1), and an enhanced scene optimization with a local neighborhood loss on the 2D grid (Section 3.2). The sorting step initializes the optimization process by explicitly organizing the values, establishing an initial smooth local neighborhood without requiring a global optimization. Once sorted, our enhanced training incorporates a neighborhood loss, further enforcing local smoothness on the 2D grid, while the differentiable renderer of the 3D Gaussian splatting optimizes the Gaussians attributes to accurately represent the scene. Similar to the original 3DGS approach, Gaussians are pruned from the scene at regular intervals, and new Gaussians are added in under-reconstructed parts of the scene. This step is called densification. After each densification step, we re-sort all Gaussians into the 2D grid. This step re-establishes local neighborhoods after pruning and densification may have disrupted them. The attribute channels cannot be sorted individually, as each position on the grid belongs to the same Gaussian splat object, as visualized in figure 3. Therefore, the sorting algorithm has to find one permutation that satisfies all attributes at once. Otherwise, it would require additional storage size to align the attributes of the sorted grids.
Figure 3: Visualization of the conversion of 2D attribute grids into Gaussian splat objects. Note that the sorted grids have to align. Therefore, they cannot be sorted individually. Here, color represents the DC components of the spherical harmonics.
To store the Gaussian attributes, we leverage off-the-shelf compression methods (described in Section3.4), which can be applied effectively to the well-organized data.
@@ 3.1 Sorting High Dimensional Gaussians into 2D Grids
Traditional grid-sorting algorithms are suitable for thousands of items, yet fall short for our application, where even small scenes contain millions of Gaussians across multiple dimensions. To address this, we combine and enhance concepts from existing sorting methods discussed in Section 2.2. Our aim is to handle the increased data volume and complexity efficiently, while ensuring that regular sorting of Gaussian data during training does not negatively affect training time. We call our newly developed algorithm Parallel Linear Assignment Sorting (PLAS).
Our strategy involves an approximation approach to the highly complex assignment problem, inspired by the principles of the Fast Linear Assignment Sorting (FLAS) algorithm [2]. We aim to find an optimal balance between computational efficiency and maintaining the accuracy of the spatial relationships among the Gaussian attributes. The parallel assignment process is visualized with random RGB colors in Figure 4: We initialize our grids by mapping the Gaussians to random positions to avoid getting stuck in a local minimum. Subsequently, we perform a low-pass filtering operation on the grid to construct an idealized target grid, and re-assign all elements to their best-matching positions regarding the smooth target. This process is repeated while decreasing the filter size, gradually sorting the grid.
@@
Organizing Gaussians into a 2D Grid: The presented algorithm is implemented highly parallel on the GPU, and will organize millions of Gaussians in a few seconds. In our application, we normalize all attributes of the Gaussians (coordinates, scale, etc.) independently to the range $[0,1]$ and then scale them with individual weights before calculating the distance of the possible re-assignments. We use the activated values for all attributes (e.g. exponentially activated scale), as they would be seen by the renderer. At the start of the training, we sort once, and continue sorting with each densification step. As the number of Gaussians grows over time, so will the the size of the grid. We find the largest square grid size that can be completely filled with Gaussians for simplicity, and prune the Gaussians with the lowest opacity values that will not fit.
@@ 3.2 Smoothness Regularization
For the regularization, we first create a blurred and thus smoother version of the 2D grid by applying a 2D Gaussian filter on all attribute channels. Deviations of the current Gaussian parameters at an original 2D grid position from those of the smoothed grid form the additional smoothness loss for training. For comparison, we select a Huber loss, as it is less sensitive to outliers than MSE. This has shown to improve the compression rate. We add the loss of the smoothness regularization to the loss of the 3D Gaussian Splatting algorithm using a weight $\lambda$. As we propagate the gradient of the smoothness regularization through the Gaussian renderer during training, we specifically enforce the 3D Gaussian Splatting optimization process to prefer Gaussians that improve the quality of the rendering while also considering the respective local smoothness. This differentiates our method from compression methods that focus on reducing existing data in a pure post-processing step. Instead, we manipulate the data during creation to achieve high compression and high visual quality afterwards.
3.3 Coordinate space contraction
We introduce an additional exponential activation for the coordinates $x$ used during training, allowing for more relative precision in coordinates in the center of the scene:
$x=\text{sign}(x_{\text{log}})\times(\exp(|x_{\text{log}}|)-1)$ (1)
Some NeRF implementations have used space contraction for unbounded scenes, notably Mip-NeRF 360 [1]. We have chosen to use this simpler logarithmic contraction function, which is easily invertible, and in design closer to the activation functions for scale and opacity used in 3DGS. It supports unbounded values, both negative and positive, as well as 0, all of which are required for the cartesian coordinates.
3.4 Quantization & Compression
Once training has finished, we can use off-the-shelf image compression methods to store the data on disk. We store the un-activated Gaussian parameters, retaining the activation methods from 3DGS: sigmoid actiation for opacity and exponential activation for the scale, and applying our own logarithmic space contraction method (see Section 3.3).
We clip the RGB values (DC values of the spherical harmonics) to the range $[-2,4]$. Opacity values are clipped to the range $[-6,12]$, non-DC SH features to $[-1,1]$, rotations to $[-1,2]$. These ranges are chosen to cover the 1st to 99th percentile of these values across all models from the 360 dataset. All but the scaling values are then normalized to the range $[0,1]$.
We quantize the different attributes by rounding them to the closest value of a linear range of $q$ total values, with $q_{\text{coords}}=2^{14}$, $q_{\text{scale}}=q_{\text{opacity}}=q_{\text{rotation}}=2^{6}$ and $q_{\text{SH\_rest}}=2^{5}$. We compress the RGB grid with lossy JPEG XL, as an 8-Bit image with a quality level of 100. All other attributes are stored as lossless JPEG XL. Our method is not restricted to this particular codec, but could also be used with other, existing 2D coding techniques.
4 Evaluation
In the following, we will perform several experiments to compare the quality and storage size to state-of-the-art 3D scene representation methods. Our experiments are based off the official 3DGS implementation on GitHub [^1] https://github.com/graphdeco-inria/gaussian-splatting. The parallel sorting is implemented in PyTorch and also uses the CUDA backend. An evaluation on the runtime performance of the sorting algorithm is given in the supplementary material. For our comparisons to prior methods, we select three real-world 3D scene reconstruction datasets: Mip-NeRF360, Tanks&Temples and Deep Blending and a synthetic dataset: Synthetic-NeRF.
@@ 4 Evaluation
Table 1: A comparison of our method to the default 3DGS with and without spherical harmonics (SH) and prior NeRF-based renderer. All sizes are in MB. Results with a $\dagger$ are directly copied from [15, 21]. Results for SSIM are included in the supplementary material.
Quantitative results: In Table 1, we compare our method to the default 3D Gaussian Splatting algorithm and prior NeRF-based 3D reconstruction methods, reporting the standard PSNR and L-PIPS metrics. Variants marked w/o SH are using only the DC part of the spherical harmonics to provide color on the Gaussian Splats, and deactivating any view-dependent effects (higher-level spherical harmonics).
We demonstrate that we reduce the average storage size by a factor of 17x to 42x (depending on the dataset), compared with 3DGS [15], without sacrificing visual quality. The highest reduction of 41.6x is observed on the Deep Blending dataset. Training our method with spherical harmonics deactivated yields a reduction factor of 127x over vanilla 3DGS while improving in PSNR.
@@
For some datasets, compared to Mip-NeRF360 [1] and VQ-TensoRF [21], we achieve a slightly higher PSNR and slightly lower L-PIPS, but at the same time allow for real time rendering and fast training, given the efficient rasterization algorithm of 3DGS [15]. Additionally, while Mip-NeRF360 trains one scene for several hours, our method only uses 10 to 30 minutes. Most notably, we achieve the same training time as 3DGS, even though our method brings some computational overhead compared to the default 3DGS, arising from the periodic sorting of Gaussians and the necessity to blur grids for computing the neighbor loss. This is due to our parameter configuration enforcing the optimization process to create a considerably smaller number of Gaussians. For example, for the Truck dataset, our method creates 1.55M Gaussians, whereas the default 3DGS creates 2.58M. For the same reason of the reduced number of Gaussians, our scenes render much faster in the 3DGS viewer than the original models. The vanilla Truck with 2.58M Gaussians renders at 385 fps with the default settings on the same GPU, while our scene with 1.55M Gaussians renders at 515 fps with better visual quality.
Figure 6: A qualitative comparison across example scenes from all four datasets.
@@
Qualitative results: The results of the metrics suggest that the rendering quality of our method is very similar to the vanilla 3DGS. This is also demonstrated in figure 6, where we directly compare the renderings from different datasets side by side. Across all examples we observe very similar quality compared to vanilla 3DGS. This underlines that our method is able to maintain a high rendering quality, while only using a fraction of the storage size.
@@
Figure 7 displays examples of the sorted 2D grids during the training process. Specifically, we show the 2D RGB color grids for thekitchen dataset. We observe that the colors of the scenes organize themselves into clusters of similar colors. The grids, however, do not appear to be more organized with more iterations. This has two reasons. Firstly, the number of Gaussians grows substantially during the training, posing an increasingly more difficult sorting task. And secondly, as our method has to sort all attributes at once using the same permutation, isolated features, such as the color cannot be organized perfectly.
@@
Figure 7: Resulting 2D attribute grids for the color. Until iteration 15000, the number of Gaussians grow from about 100k up to 1.5 million.